OpenStudy (kammykoo):

What number should be added to both sides of the equation to complete the square? x2 – 6x = 5 What number should be added to both sides of the equation to complete the square? x2 – 6x = 5

OpenStudy (sachintha):

\(\sf\underline{Completing~squares}\) \(\large ax^2+bx=0\) Divide the whole equation by the coefficient of \(x^2\). \(\large x^2+\Large\frac{b}{a}\large x=0\) Divide the coefficient of \(x\) by 2. \(\large x^2+\Large\frac{b}{2a}\large x=0\) Add the squared new coefficient of \(x\) to both sides. \(\large x^2+\Large\frac{b}{2a}\large x+(\Large\frac{b}{2a})\large^2=0+\Large(\frac{b}{2a})\large^2\) \(\large(x+\Large\frac{b}{2a})\large^2=0+\Large(\frac{b}{2a})\large^2\)